OpenStudy (anonymous):

find antiderivative of 1/(5th root of t^4)

OpenStudy (amistre64):

1 --------- ? 5root(t^4)

OpenStudy (amistre64):

1/t^(4)^(1/5) = 1/t^(4/5) = t^(-4/5)

OpenStudy (amistre64):

t^(-4/5 + 5/5) ------------ -4/5 + 5/5

OpenStudy (amistre64):

5 * 5root(t)

OpenStudy (amistre64):

.... + C :)